Question 9 • EHS Content Recovery - Probability and Statistics - Part C
The sampling distribution of S − F has mean 0.31 − 0.38 = −0.07 and standard deviation √[(0.31)(0.69)/45 + (0.38)(0.62)/40] ≈ 0.103. The z-score for S − F = 0 is [0 − (−0.07)]/0.103 ≈ 0.679. Thus, P(S > F) = P(Z > 0.679) ≈ 1 − 0.752 = 0.248, choice A.
AI-written and checked against the verified answer.